OpenStudy (anonymous):

(3m^2n)(-6m^3)(2/9m^2)=?

OpenStudy (callisto):

Is the question \((3m^2n)(-6m^3)(\frac{2}{9}m^2)\) ?

OpenStudy (anonymous):

yes :)

OpenStudy (nurali):

\[(3m^2n)*(-6m^3)(\frac{ 2 }{ 9 }m^2)\] \[(\frac{ -36m^7n }{ 9 })\] \[-4m^7n\]